OpenStudy (anonymous):

Whats the focus of a parabola centered at (0,0) and using the point (4,6)? Help pleasee.

OpenStudy (anonymous):



OpenStudy (anonymous):

We have this two possibilities...anything else given for the problem?

OpenStudy (anonymous):

I know its going to be pointing up!

OpenStudy (anonymous):

A parabola doesn't have a center. I think you meant vertex (0, 0). As arghya stated, there are two parabola that pass the point (4, 6) and thus two foci. If you want the one facing up, then x^2 = 4py 4^2 = 4p(6) p = 16/24 = 2/3 The focus is (0, 2/3)

OpenStudy (anonymous):

and it would be different for y^2=4ax...thn it would be 9/4,0

OpenStudy (anonymous):

thank ya`ll so much (: